OpenStudy (anonymous):

If a rectangle has a perimeter of 70, a length of x and a width of x - 9, find the value of the length of the rectangle.

OpenStudy (ybarrap):

Perimeter is sum of length and width: $$ \text{length}+\text{width}=x+(x-9)=70 $$ Solve for x, then length = x.

OpenStudy (anonymous):

Thanks!

OpenStudy (mertsj):

Perimeter is sum of twice width + twice length

OpenStudy (mertsj):

2x+2x-18=70

OpenStudy (ybarrap):

That is true! Thank @Mertsj

OpenStudy (mertsj):

yw

OpenStudy (anonymous):

I got it now. Thanks!